During peak traffic periods, accidents occur in a certain small city at a rate of 4.9 per hour. The morning peak lasts for 1.6 hours. On any given day, the probability that there will be at most 2 accidents during the morning peak period is


1
Expert's answer
2021-11-11T11:26:58-0500

The probability that there will be at most 2 accidents during the morning peak period is:

"P(X\eqslantless2)=P(X=0)+P(X=1)+P(X=2)"


"P(X\eqslantless2)=\frac{1.6^{0}*e^{-1.6}}{0!}+\frac{1.6^{1}*e^{-1.6}}{1!}+\frac{1.6^{2}*e^{-1.6}}{2!}"


"P(X\eqslantless2)=0.2019+0.3230+0.2584=0.7834"


There is 78.34% probability of that there will be at most 2 accidents during the morning peak period.